The Hubble Space Telescope is in a "Low-Earth Orbit" with an orbital period of 95min. Calculate the altitude of the HST's orbit above the earth's surface. Express your answer either in meters.

Earth's radius: 6.371x106 m

Earth's mass: 5.97x1024kg
1
Expert's answer
2020-04-30T10:41:23-0400

Solution.

"T = 95 min = 5700s;"

"R_{\bigoplus} = 6.371\sdot10^6m;"

"M_{\bigoplus} = 5.97\sdot10^{24}kg;"

"H - ?;"

The linear velocity of a satellite orbiting the Earth is sought by the formula:

"\upsilon = \sqrt{G\sdot \dfrac{M_{\bigoplus}}{R_{\bigoplus}+H}};"

In addition, the linear velocity of the body in a circle is sought by the following formula:

"\upsilon = \dfrac{2\pi R}{T};"

For our case we have:

"\dfrac{2\pi( R_{\bigoplus} +H)}{T} = \sqrt{G\sdot \dfrac{M_{\bigoplus}}{R_{\bigoplus}+H}};"



"\dfrac{4\pi^2( R_{\bigoplus} +H)^2}{T^2} = {G\sdot \dfrac{M_{\bigoplus}}{R_{\bigoplus}+H}};"

"\dfrac{4\pi^2( R_{\bigoplus} +H)}{T^2} = {G\sdot M_{\bigoplus}};"

"R_{\bigoplus} +H =\dfrac{ G\sdot M_ {\bigoplus}T^2}{4\pi^2};"

"H = \dfrac{ G\sdot M_ {\bigoplus}T^2}{4\pi^2} - R_{\bigoplus};"

"H = \dfrac{6.67\sdot 10^{-11}\sdot5.97\sdot10^{24}\sdot5700^2}{4\sdot3.14^2} -6.371\sdot10^6="

"=3.28\sdot10^{20}m;"

Answer: "H = 3.28\sdot10^{20}m."